Circular No. 6842

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

η CARINAE

M. F. Corcoran, Goddard Space Flight Center (GSFC) and Universities Space Research Association; J. H. Swank and R. Petre, GSFC; and K. Ishibashi and K. Davidson, University of Minnesota, report: “Continuing observation of this peculiar luminous star with the RXTE PCA shows that the star may be emerging from the x-ray ‘low state’ that began in early Dec. 1997. During the low state, the PCA countrate in the band 2–10 keV dropped from > 70 to 15 PCU counts/s in a period of ∼ 1 month. Recent RXTE observations show that the observed x-ray emission has risen by about a factor of 2 since 1998 Feb. 2. After correcting for the permanent sky background, the countrate due to η Car itself has increased by a factor close to 6. Since Feb. 16, the average rate of increase in the x-ray emission has been ΔR = +0.7 PCU counts s^{−1} day^{−1}. This is still considerably less than the rate of decline to x-ray minimum, ΔR = −1.6 PCU counts s^{−1} day^{−1} during the interval 1997 Nov. 11–Dec. 18. Using a conversion factor of 3.22 × 10^{−12} erg s^{−1} cm^{−2} per PCA count in the band 2–10 keV suggests that at present the rate of increase in 2–10-keV flux is ∼ 2.25 × 10^{−12} erg cm^{−2} s^{−1} day^{−1}. If the rate of increase remains constant, then the x-ray flux from η Car should reach or exceed the previous maximum flux level on May 8. The recent x-ray brightening could also be associated with episodic flaring previously reported (IAUC 6668), since the next flare maximum is expected to occur on Mar. 17.”

1SAX J1603.9–7753 J. M. Muller, BeppoSAX Science Data Center, Rome, and Space Re-

search Organization of the Netherlands (SRON), Utrecht; J. in ’t Zand, SRON; M. J. S. Smith, BeppoSAX Science Operation Center (SOC), Rome, and SRON; A. Coletta, L. Di Ciolo, and V. Torroni, SOC; and L. Piro, Istituto Astrofisica Spaziale, CNR, Rome, report on behalf of the BeppoSax team: “On Mar. 6.16 UT, the BeppoSAX Wide Field Camera 2 detected an x-ray burst-like event at α = 16^{h}03^{m}54^{s}, δ = −77^{o}53′.1 (equinox 2000.0, preliminary error radius 3′). As no catalogued x-ray source can be found within the error circle, we call the source 1SAX J1603.9–7753. The event lasted ∼ 40 s and had peak intensities of 0.15 Crab in the energy band 2–10 keV and 0.30 Crab at 10–25 keV. The rise and decay times were similar. The source was not detected during 1.1 days before the burst, with a 3σ upper limit of 2 mCrab (2–10 keV). The error circle contains the Tycho star 9446-00734-1 (distance 80 pc), as well as IRAS 15567–7741.”

1998 March 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green